OpenStudy (anonymous):

Help please: x(T)= -1, t<0 0, t=0 1, t>0 system out: y(t)= u(t)(1-2(e^-t)) -2u(t)e^(-t)-u(-t) a) Fourier transform x(t) b) Fourier transform y(t)

OpenStudy (anonymous):

fourier transform of x(t) is \[\int\limits_{-\infty}^{\infty} x(t) e ^{j \omega t} \delta t\]

OpenStudy (anonymous):

so, \[\int\limits_{-\infty}^{\infty}xdx = \int\limits_{-\infty}^{0-} xdx+\int\limits_{0-}^{0+}xdx+\int\limits_{o+}^{\infty}\]

OpenStudy (anonymous):

oops the last term is\[\int\limits_{0+}^{\infty}xdx\]

OpenStudy (anonymous):

so, substituting, we get, \[\int\limits_{-\infty}^{+\infty}x(t)e ^{j \omega t} \delta t = \int\limits_{-\infty}^{0-}-1e^{j \omega t} \delta t +0+\int\limits_{0+}^{\infty}e^{j \omega t} \delta t \]

OpenStudy (anonymous):

\[=\int\limits_{0+}^{\infty} -1/e^{j \omega t} \delta t + \int\limits_{0+}^{\infty} e^{j \omega t} \delta t \]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes, thanks, I think that I understand. I'm looking for the impulse response can u help me with that?

OpenStudy (anonymous):

the second one you mean

OpenStudy (anonymous):

i'm looking at that. hold on.

OpenStudy (anonymous):

remind me again, u(t) = 1 for t = 0 and u(t) = 0 for all other values. Am i right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so y(t) = u(t)(1-2(e^-t)) -2u(t)e^(-t)-u(-t) = u(t) - 2u(t)e^-t - 2u(t)e^-t-u(-t) = u(t)-u(-t)-4u(t)e^-t

OpenStudy (anonymous):

is u(-t) = -u(t)?

OpenStudy (anonymous):

no it isn't.

OpenStudy (anonymous):

don't bother. it is not -u(t)

OpenStudy (anonymous):

so, fourier transform of y(t) is \[\int\limits_{0-}^{0+}u(t)-u(-t)-4e ^{-t} e ^{j \omega t } \delta t\]

OpenStudy (anonymous):

notice the limits

OpenStudy (anonymous):

nice! agree

OpenStudy (anonymous):

now, u(t) = 1 for t = 0+ or t = 0- so the u(t) terms cancel out. (i am not sure about this. ask some one else about this)

OpenStudy (anonymous):

what did you get as the fourier transform of y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

impulse response and input x output y , how relate

OpenStudy (anonymous):

impulse response is the response of a function at t = 0

OpenStudy (anonymous):

or rather, when a brief signal, or impulse is input, the output that you get from the system is called impulse response

OpenStudy (anonymous):

ok directly from the previous limit

OpenStudy (anonymous):

i am not sure, but I think the fourier transform of the impulse response is 0.

OpenStudy (anonymous):

One more question please: If we enter another signal to the system and

OpenStudy (anonymous):

Y{2t}=h(t)*X2{t}

OpenStudy (anonymous):

\[\sum_{-\infty}^{+\infty} C1e ^{jl2t} * h(t)

OpenStudy (anonymous):

\[\sum_{-\infty}^{+\infty} C1e ^{jl2t} * h(t) \]

OpenStudy (anonymous):

i am sorry. Post it again tomorrow. I am off to bed :)

OpenStudy (anonymous):

ok dont problem thanks 4 erything